OpenStudy (anonymous):

Part 1: Find the solution to the equation 64(2 � x) = 44x. (4 points) Part 2: Using complete sentences, explain the procedure used to solve this equation. (4 points)

OpenStudy (anonymous):

cut and paste fail

OpenStudy (anonymous):

its 64(2-x)=44x

OpenStudy (anonymous):

Just kidding its 64^(2-x)=4^4x

OpenStudy (anonymous):

Did you ever find the answer?